# Foundation Models Meet Low-Cost Sensors: Test-Time Adaptation for Rescaling Disparity for Zero-Shot Metric Depth Estimation

### v1

@@ Abstract
The recent development of foundation models for monocular depth estimation such as Depth Anything paved the way to zero-shot monocular depth estimation. Since it returns an affine-invariant disparity map, the favored technique to recover the metric depth consists in fine-tuning the model. However, this stage is costly to perform because of the training but also due to the creation of the dataset. It must contain images captured by the camera that will be used at test time and the corresponding ground truth. Moreover, the fine-tuning may also degrade the generalizing capacity of the original model. Instead, we propose in this paper a new method to rescale Depth Anything predictions using 3D points provided by low-cost sensors or techniques such as low-resolution LiDAR, stereo camera, structure-from-motion where poses are given by an IMU. Thus, this approach avoids fine-tuning and preserves the generalizing power of the original depth estimation model while being robust to the noise of the sensor or of the depth model. Our experiments highlight improvements relative to other metric depth estimation methods and competitive results compared to fine-tuned approaches. Code available at gitlab.ensta.fr/ssh/monocular-depth-rescaling.
@@ I Introduction
Despite the growth of 3D perception sensors such as LiDAR, time-of-flight, structured light or stereo cameras, for robotic systems, traditional monocular cameras remain a key sensor for any robot setup. In addition to being a cheaper solution, monocular depth estimation can also offer denser outputs as well as larger depth range. The increase in the number of open datasets for monocular depth estimation and the development of neural network architectures such as Vision Transformers that scale well with the size of the training dataset [1] allowed the emergence of foundation models for monocular depth estimation [2].
@@
However, predicting the depth from a single image is an ambiguous task since depth depends on the image content and on the camera calibration. The consequence is that models trained on a single dataset to make metric depth (also referred to as absolute depth) predictions generalize poorly on images captured with a different camera. To avoid this issue during training on multiple datasets, some methods [3, 2, 4] learn an affine-invariant depth or disparity which allows impressive results on zero-shot relative depth estimation benchmarks. The most common way of adapting these models in order to make metric depth predictions consists in fine-tuning on the target domain, i.e., on a dataset composed of images captured with the target camera calibration. Such a solution is costly due to the creation of the dataset and the training computation. Moreover, transferring performance on images with a different calibration still remains impossible. Several solutions have been proposed to solve this issue by taking into account the camera calibration in the method. However, such approaches are often more costly at inference or cannot be trained on image datasets with unknown calibration such as ImageNet [5] contrary to methods like Depth Anything [2, 4].
@@
In this paper, we investigate the use of low-cost sensors for solving the scale ambiguity in order to perform zero-shot metric depth estimation given affine-invariant disparity predictions. The additional sensor provides some reference 3D points that are exploited to recover the scaling parameters. We note for the sake of brevity, we will refer to rescaling the process of finding an affine transformation and applying it to recover the metric depth. We explored several low-cost sensors including low-resolution LiDARs (with 16 and 32 beams) as well as 2D LiDARs (with a single beam) that are often used for indoor robotics, stereo cameras and IMU. The latter provides the relative camera pose that is leveraged to triangulate matching points in successive images in a structure-from-motion strategy. The advantage of such an approach is twofold. On the one hand, it can be used with any monocular depth estimation model such as Depth Anything V1 and V2 [2, 4] with a high generalization ability due to their large and diverse training dataset. On the other hand, our method does not require any costly fine-tuning on the target domain and provides instant adaptation. We conducted extensive experiments that demonstrate the interest of our approaches on standard metric depth estimation benchmarks.
@@ II Related work
Figure 1: Illustration of our method. First, an affine-invariant disparity map $d$ is predicted from an image with a neural network such as Depth Anything [2] while in parallel a low-cost sensor is used to estimate a set of 3D points $P$. $P$ and the corresponding values in $d$ are then used to estimate the scaling parameters $\alpha$ and $\beta$ using a RANSAC [6]. The parameters are applied to $d$ to recover the metric depth $D$.
@@ II-A Monocular depth estimation
While pioneer works on monocular depth estimation relied on Markov Random Fields [7, 8], subsequent ones have shown the effectiveness of convolutional neural networks [9, 10] then transformers for this task [11]. More recently, benefiting from advances in image generation [12], impressive results have been obtained using diffusion models [13, 14]. Regarding the output, monocular depth estimation has been initially addressed as a regression problem [9], before moving to classification approaches with discrete bins that showed better performances [15, 16, 17]. Monocular depth estimation methods can also be divided whether they are supervised [9, 10, 15] or self-supervised [18]. In the latter case, depth is learned with a photometric loss that compares a target image with its reconstruction obtained from another view of the same scene, the depth prediction and the camera pose. Recently, the multiplication of depth estimation benchmark allowed training models on multiple domains at once [3] and paved the way to zero-shot monocular depth estimation. Due to the inherent scale ambiguity of depth estimation from a single image, these methods are mostly trained to produce affine-invariant depth or disparity predictions [3, 11, 2]. Our work aims to be used with any depth estimation model as long as it returns disparity maps that are accurate within an affine transformation whatever its other characteristics in terms of training set or architecture.
@@ II-B Scale estimation for monocular depth
A major issue of monocular depth prediction relies on scale ambiguity which means that the true size of an object cannot be recovered for sure from a single image. Since most monocular depth estimation models are trained for a specific camera calibration, using them with another camera leads to ill-scaled predictions. Likewise, self-supervised depth estimation methods may also predict depth up to an affine transformation if they are trained with no reference of absolute distance at all (most of the time when the camera pose is also estimated by a neural network). The most common way to recover the metric depth in this situation is to fine-tune the model on a dataset collected with the camera that will be used at inference. However, this solution is costly to implement because it needs the creation of an image dataset with the relative ground truth and a new training. In case of self-supervised depth estimation, the scaling factor can be recovered from the ground truth pose given by an IMU to compute the photometric loss [19], the constraints of a stereo camera [20, 18, 21] or the speed of the vehicle that carries the camera [22, 23]. Other works focus on the temporal consistency of the scale of depth predictions [24, 25]. Closer to our work, [26] learns to predict scaled depth maps from affine-invariant disparity maps and a visual-inertial odometry. In [27], authors propose to estimate the scale factor for a target domain from a model that has been jointly trained in a self-supervised manner on a source domain with known ground truth and on a target domain without depth annotation. In contrast, we propose to rescale any disparity prediction that is correct up to an affine transformation with no training or fine-tuning at all but by leveraging reference 3D points provided by an external low-cost sensor. Furthermore, the use of an external sensor makes our approach adaptable to any camera calibration.
II-C Monocular zero-shot metric depth estimation
@@
ZoeDepth [28] is the first zero-shot metric depth estimation method. It first consists in relative depth pre-training of the MiDaS [3] backbone then fine-tuning two metric bins modules, one for indoor scenes and the other for outdoor scenes. More recently, several zero-shot metric depth estimation methods have been proposed without any fine-tuning. Thus, ScaleDepth [29] decomposes metric depth estimation in relative depth estimation and scale estimation each with a dedicated module. It can also leverage a text description of the scene to guide the supervision. In [30], authors introduce a dedicated architecture that takes as input the calibration matrix in addition to the image. On the contrary, [31] makes predictions from the images only but applies transformation on the input images so the predictions are invariant to the image size or the camera calibration. UniDepth [32] estimates an internal representation of the camera calibration directly from the input images only. These approaches still have drawbacks since they are often more costly at inference or need for the calibration of the images even at the training stage. This prevents exploiting image datasets with unknown calibration as Depth Anything V1 [2] and V2 [4] which leverage a distillation strategy with such datasets to improve their generalization abilities.
@@ III Method
Let $\phi$ be a monocular depth estimation model such as [3, 2] that is trained to predict an affine-invariant disparity map $d\in\mathbb{R}^{H\times W}$ given an input image $I\in\mathbb{R}^{H\times W\times 3}$ where $H$ and $W$ are the height and the width of the image $I$. Therefore, the metric or absolute depth map $D\in\mathbb{R}^{H\times W}$ that corresponds to the inverse of the metric or absolute disparity $D^{-1}$ given by the relation:
@@
Our method, illustrated in Fig. 1, aims at recovering the metric depth map $D$ from the affine-invariant disparity map $d$ and a set of $N$ reference 3D points $P\in\mathbb{R}^{N\times 3}$ by regressing the parameters $\alpha$ and $\beta$. First, we perform a bilinear sampling on the affine-invariant map $d$ at the locations of the projection of the reference 3D points $P$ on the image plan so as to have an affine-invariant disparity value for each reference 3D point. Second, we leverage a linear regression to estimate parameters $\alpha$ and $\beta$. To favor the robustness to potential outliers in the reference 3D points, we use a RANSAC algorithm [6], but it could be replaced by another robust regression method.
In our approach, we assume the set $P$ is provided with a low-cost sensor like a low-resolution LiDAR, a stereo camera or an IMU from which the poses of the camera that capture $I$ can be extracted. The poses are used in a structure-from-motion (SFM) strategy to triangulate matching points in consecutive images of a video. Thus, $P$ may be a very sparse depth map, i.e., $N\ll HW$ especially when provided by a low-resolution LiDAR or by a structure-from-motion technique. It may also contain some outlier measures for instance the depth of the most distant pixel given by a stereo camera or when applying structure-from-motion with dynamic objects.
As our method relies on affine-invariant disparity maps, we can exploit, without any fine-tuning, foundation models like Depth Anything V1 or V2 [2, 4] that are trained to predict such outputs. Furthermore, the robustness to outliers enabled by the RANSAC allows low-quality sensors to be used to obtain the reference 3D points. Thus, we can provide low-cost metric depth maps whatever the environment or the camera. Nevertheless, since Depth Anything processes each image independently and is supervised in such a way there is no guarantee its disparity predictions are normalized identically even for consecutive images of a video, the scaling factor $\alpha$ and the offset $\beta$ must be estimated for each image.
@@ IV-A Dataset and metrics
To evaluate the performance in zero-shot metric depth estimation, we focus on standard benchmarks for monocular depth estimation that have not been used to pretrain Depth Anything [2]. This includes indoor datasets: NYU V2 [33], SUN-RGBD [34], IBIMS-1 [35] and DIODE indoor [36] and outdoor datasets: KITTI [37] , DDAD [23] and DIODE outdoor [36]. We also use DIML dataset [38] that appears in the training set of Depth Anything to show the limitations of rescaling with stereo depth. We adopt standard depth estimation metrics (see [9]) to compare our method to other approaches of the literature.
@@ IV-B Implementation details
All experiments have been conducted using Depth Anything V1 [2] with ViT Large [1] without fine-tuning unless otherwise mentioned. We adopt Depth Anything V1 [2] code base and settings for evaluation.
We simulate low-resolution LiDARs by evenly selecting as many horizontal lines in the ground truth depth maps as the number of beams in 32-laser, 16-laser or 2D LiDARs (i.e., with a single beam). We study rescaling using stereo depth on KITTI dataset only since it provides stereo images. For this purpose, we estimate stereo depth maps using Semi Global Matching (SGM) [39] with the settings from [40]. We study the structure from motion for KITTI [37] and DDAD datasets [23] as they both consist of temporal sequences of images including the pose estimations. To obtain the reference 3D points, we first extract matching keypoints in the target image and the previous one using SIFT [41] or OmniGlue [42] and triangulate them using the pose between these images. Since this strategy requires enough displacement magnitude between the two images, we only consider image couples with a rotation higher than $5$ degrees and a translation greater than $1.5$ and $2$ meters for KITTI and DDAD datasets, respectively. We note that this threshold is the only hyperparameter that needs to be tunned in our approach.
IV-C Results
TABLE I: Quantitative results on the NYU V2 benchmark [33] (indoor).
Methods $\delta_{1}$uparrow $\delta_{2}$uparrow $\delta_{3}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ log10 $\downarrow$
ZeroDepth [30] (ZS) 0.926 0.986 0.081 0.338
Metric3D (ZS) 0.944 0.986 0.995 0.083 0.310 0.035
Metric3D V2 [31] (ZS) 0.975 0.994 0.998 0.063 0.251 0.028
@@ IV-C Results
[…]
Ours rescaling w/ LiDAR 1 beam 0.939 0.981 0.993 0.063 0.652 0.028
[…]
TABLE II: Quantitative results on the KITTI benchmark [37] (outdoor).
Methods $\delta_{1}$uparrow $\delta_{2}$uparrow $\delta_{3}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ RMSE Log $\downarrow$
Semi-Global Matching [39] (stereo) 0.855 0.947 0.976 0.160 7.351 0.227
ZeroDepth [30] (ZS) 0.910 0.980 0.996 0.102 4.044 0.172
Metric3D (ZS) 0.964 0.993 0.998 0.058 2.770 0.092
[…]
ScaleDepth [29] (FT) 0,980 0,980 1.000 0,048 1,987 0,073
[…]
Ours rescaling w/ LiDAR 1 beam 0.891 0.974 0.986 0.131 3.096 0.142
[…]
Ours rescaling w/ SFM (SIFT [41]) 0.893 0.985 0.998 0.103 3.920 0.134
[…]
Ours rescaling w/ stereo (SGM [39]) 0.941 0.989 0.997 0.072 3.526 0.111
TABLE III: Quantitative results on different zero-shot indoor benchmarks.
Methods SUN-RGBD IBIMS-1 DIODE Indoor
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
Metric3D 0.144 0.252
@@
TABLE IV: Quantitative results on different zero-shot outdoor benchmarks.
Methods DIODE Outdoor DDAD
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
ZoeDepth [28] (FT KITTI) 0.814 0.156 10.678
@@
Depth Anything [2] (FT KITTI) 0.288 0.794 6.641 0.886 0.105 5.931
Ours rescaling w/ Lidar 1 beam 0.689 0.880 6.222 0.706 0.326 9.229

### v2

# A Simple yet Effective Test-Time Adaptation for Zero-Shot Monocular Metric Depth Estimation

@@ Abstract
The recent development of foundation models for monocular depth estimation such as Depth Anything paved the way to zero-shot monocular depth estimation. Since it returns an affine-invariant disparity map, the favored technique to recover the metric depth consists in fine-tuning the model. However, this stage is not straightforward, it can be costly and time-consuming because of the training and the creation of the dataset. The latter must contain images captured by the camera that will be used at test time and the corresponding ground truth. Moreover, the fine-tuning may also degrade the generalizing capacity of the original model. Instead, we propose in this paper a new method to rescale Depth Anything predictions using 3D points provided by sensors or techniques such as low-resolution LiDAR or structure-from-motion with poses given by an IMU. This approach avoids fine-tuning and preserves the generalizing power of the original depth estimation model while being robust to the noise of the sparse depth or of the depth model. Our experiments highlight enhancements relative to zero-shot monocular metric depth estimation methods, competitive results compared to fine-tuned approaches and a better robustness than depth completion approaches. Code available at gitlab.ensta.fr/ssh/monocular-depth-rescaling.
@@ I Introduction
Despite the growth of 3D perception sensors such as LiDAR, time-of-flight, structured light or stereo cameras, for robotic systems, traditional monocular cameras remain a key sensor for any robot setup. In addition to being a cheaper solution, monocular depth estimation can also offer denser outputs as well as a larger depth range. The increase in the number of open datasets for monocular depth estimation and the development of neural network architectures such as Vision Transformers that scale well with the size of the training dataset [1] allowed the emergence of foundation models for monocular depth estimation [2].
@@
Predicting metric depth (also referred to as absolute depth) from a single image is fundamentally an impossible task due to scaling ambiguities. However, monocular depth estimation methods [3] can achieve very good performances in a defined context. These approaches have learned depth cues in images sampled from a certain distribution of environments that have been captured by a camera with fixed calibration. Thus, such models, trained on a single dataset, generalize poorly to images taken with a different camera. To avoid this issue during training on multiple datasets, some methods [4, 2, 5] learn an affine-invariant depth or disparity which allows impressive results on zero-shot relative depth estimation benchmarks. Then, they propose to fine-tune their pre-trained models on the target domain to make metric depth predictions, i.e., on a dataset composed of images captured with the target camera calibration. Such a solution is costly in real cases due to the creation of the dataset and the training computation. Moreover, it must be performed for each new camera calibration. Several solutions have been proposed to solve this issue by explicitly taking into account the camera calibration in the method [6] or trying to learn it [7]. Depth completion methods propose another alternative [8, 9] as they take as input some sparse depth measurements. However, all these approaches may be more costly at inference or cannot be trained on image datasets with unknown calibration or unknown ground truth depth such as ImageNet [10] contrary to methods like Depth Anything [2, 5].
@@
In this paper, following depth completion approaches, we investigate a test-time adaptation that leverages sparse depth measurements for solving the scale ambiguity in order to perform zero-shot monocular metric depth estimation given affine-invariant disparity predictions. Thus, an additional sensor is used to obtain some reference 3D points that are exploited to recover the scaling parameters. For the sake of brevity, we will refer to rescaling the process of finding an affine transformation and applying it to recover the metric depth. We focus our study on sparse depth that can be provided by other sources including low-resolution LiDARs (with 16 and 32 beams), 2D LiDARs (with a single beam) that are often used for indoor robotics and structure-from-motion. In the latter, we assume a metric relative camera pose is given by an IMU. The advantage of our approach is twofold. On the one hand, it can be used with any monocular depth estimation model such as Depth Anything V1 and V2 [2, 5] with a high generalization ability due to their large and diverse training dataset. On the other hand, our method does not require any costly fine-tuning on the target domain and provides instant adaptation. We conducted extensive experiments to evaluate our approaches on standard metric depth estimation benchmarks and demonstrate robustness to a noisy sparse depth or to a drop in sparse depth density.
@@ II Related work
Figure 1: Illustration of our method. First, an affine-invariant disparity map $d$ is predicted from an image with a neural network such as Depth Anything [2] while in parallel a sensor is used to estimate a set of 3D points $P$. $P$ and the corresponding values in $d$ are then used to estimate the scaling parameters $\alpha$ and $\beta$ using a RANSAC [11]. The parameters are applied to $d$ to recover the metric depth $D$.
@@ II-A Monocular depth estimation
While pioneer works on monocular depth estimation relied on Markov Random Fields [12, 13], subsequent ones have shown the effectiveness of convolutional neural networks [14, 15] then transformers for this task [16]. More recently, benefiting from advances in image generation [17], impressive results have been obtained using diffusion models [18, 19]. Regarding the output, monocular depth estimation has been initially addressed as a regression problem [14], before moving to classification approaches with discrete bins that show better performances [3, 20]. Monocular depth estimation methods can also be divided into whether they are supervised [14, 15, 3] or self-supervised [21]. Recently, the multiplication of depth estimation benchmark allowed training models on multiple domains at once [4] and paved the way to zero-shot monocular depth estimation. Due to the inherent scale ambiguity of depth estimation from a single image, these methods are mostly trained to produce affine-invariant depth or disparity predictions [4, 16, 2]. Our work aims to be used with any depth estimation model as long as it returns disparity maps that are accurate within an affine transformation whatever its other characteristics in terms of training set or architecture.
@@ II-B Scale estimation for monocular depth
A major issue of monocular depth prediction relies on scale ambiguity which means that the true size of an object cannot be recovered for sure from a single image. Since most monocular depth estimation models are trained for a specific camera calibration, using them with another camera leads to ill-scaled predictions. The most common way to recover the metric depth in this situation is to fine-tune the model on a dataset collected with the camera that will be used at inference [4, 16]. In practice, this solution is costly to implement as it requires the creation of an image dataset with the relative ground truth and a new training. Other works focus on the temporal consistency of the scale of depth predictions [22, 23]. Closer to our work, [24] learns to predict scaled depth maps from affine-invariant disparity maps and visual-inertial odometry. Also, there exists an extensive literature on depth completion which studies neural network architectures that take as inputs both an image and a sparse depth map [25, 8, 9]. In [26], authors propose to estimate the scale factor for a target domain from a model that has been jointly trained on a source domain with known ground truth and on a target domain without depth annotation. In contrast, we propose to rescale at test time any disparity prediction that is correct up to an affine transformation with no additional training or fine-tuning but by leveraging reference 3D points provided by an external sensor or technique. Furthermore, relying on external sensor makes our approach adaptable to any camera calibration.
II-C Zero-shot monocular metric depth estimation
@@
ZoeDepth [27] is the first zero-shot monocular metric depth estimation method. It first consists in relative depth pre-training of the MiDaS [4] backbone then fine-tuning two metric bin modules, one for indoor scenes and the other for outdoor scenes. More recently, other methods have been proposed without any fine-tuning. Thus, ScaleDepth [28] decomposes metric depth estimation in relative depth estimation and scale estimation each with a dedicated module. It can also leverage a text description of the scene to guide the supervision. In [6], authors introduce a dedicated architecture that takes as input the calibration matrix in addition to the image. On the contrary, [29, 30] make predictions from the images only but applies transformation on the input images so the predictions are invariant to the image size or the camera calibration. UniDepth [7] estimates an internal representation of the camera calibration directly from the input images only. These approaches still have drawbacks since they are often more costly at inference or need the calibration of the images even at the training stage. This prevents exploiting image datasets with unknown calibration unlike Depth Anything V1 [2] and V2 [5] which leveraged a distillation strategy with such datasets to improve their generalization abilities.
@@ III Method
Let $\phi$ be a monocular depth estimation model such as [4, 2] that is trained to predict an affine-invariant disparity map $d\in\mathbb{R}^{H\times W}$ given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$ where $H$ and $W$ are the height and the width of the image $I$. Therefore, the metric or absolute depth map $D\in\mathbb{R}^{H\times W}$ that corresponds to the inverse of the metric or absolute disparity $D^{-1}$ is given by the relation:
@@
Our method, illustrated in Fig. 1, aims at recovering the metric depth map $D$ at test time from the affine-invariant disparity map $d$ and a set of $N$ reference 3D points $P\in\mathbb{R}^{N\times 3}$ by regressing the parameters $\alpha$ and $\beta$. First, we perform a bilinear sampling on the affine-invariant map $d$ at the locations of the projection of the reference 3D points $P$ on the image plan so as to have an affine-invariant disparity value for each reference 3D point. Second, we leverage linear regression to estimate $\alpha$ and $\beta$ parameters. To favor the robustness against potential outliers in the reference 3D points, we use a RANSAC algorithm [11], but other robust regression methods could be considered.
In our approach, we assume the set $P$ is provided by a low-resolution LiDAR or a structure-from-motion (SFM) technique in which poses are given by an IMU. When SFM is leveraged, having metric poses is necessary to triangulate the absolute coordinates of matching points in consecutive video images. Thus, $P$ may be a very sparse depth map, i.e., $N\ll HW$ and may also contain some outlier measurements such as those caused by LiDAR reflections from dust or artifacts introduced by SFM in dynamic scenes.
As our method relies on affine-invariant disparity maps, we can exploit, without any fine-tuning, foundation models like Depth Anything V1 or V2 [2, 5] that are trained to predict such outputs. Furthermore, the robustness to outliers enabled by the RANSAC allows low-quality sensors to be used to obtain the reference 3D points. Thus, we can provide metric depth maps whatever the environment or the camera with no fine-tuning. Nevertheless, since Depth Anything processes each image independently and is supervised in such a way there is no guarantee its disparity predictions are normalized identically even for consecutive images of a video, the scaling factor $\alpha$ and the offset $\beta$ must be estimated for each image.
@@ IV-A Dataset and metrics
To evaluate performance in zero-shot monocular metric depth estimation, we focus on standard monocular depth estimation benchmarks that have not been used to pretrain Depth Anything [2]. This includes indoor datasets: NYUv2 [31], SUN-RGBD [32], IBIMS-1 [33] and DIODE indoor [34] and outdoor datasets: KITTI [35] , DDAD [36] and DIODE outdoor [34]. We adopt standard depth estimation metrics (see [14, 2]) to compare our method to other approaches of the literature:
$\text{RMS}=\sqrt{\frac{1}{|\Omega|}\sum_{p\in\Omega}(\hat{D}(p)-D^{*}(p))^{2}}$ (2)
$\text{AbsRel}=\frac{1}{|\Omega|}\sum_{p\in\Omega}\frac{|\hat{D}(p)-D^{*}(p)|}{D^{*}(p)}$ (3)
$\delta_{1}=\frac{1}{|\Omega|}\Bigg{|}\Bigg{\{}p\in\Omega\bigg{|}\frac{1}{1.25}<\frac{\hat{D}(p)}{D^{*}(p)}<\frac{1.25}{1}\Bigg{\}}\Bigg{|}$ (4)
where $\Omega$ is the set of pixels for which the ground truth is available and |.||.|| . | applied to a set returns its cardinal and the absolute value of a scalar otherwise. For a pixel $p\in\Omega$, $\hat{D}(p)$ is the estimation corresponding the ground truth depth $D^{*}(p)$.
@@ IV-B Implementation details
All experiments have been conducted using Depth Anything V1 [2] with ViT Large [1] without fine-tuning unless otherwise mentioned. We adopt Depth Anything V1 [2] code base and settings for evaluation except for comparison with depth completion methods as detailed later. We simulate low-resolution LiDARs by evenly selecting as many horizontal lines in the ground truth depth maps as the number of beams in 32-laser, 16-laser or 2D LiDARs (i.e., with a single beam). In contrast, no ground truth depth is used as input when rescaling with SFM. We study rescaling with SFM for KITTI [35] and DDAD datasets [36] only as they both consist of temporal sequences of images including the pose estimations. To obtain the reference 3D points, we first extract matching keypoints in the target image and the previous one using SIFT [37] or OmniGlue [38] and triangulate them using the pose between these images. Since this strategy requires enough displacement magnitude between the two images, we only consider image couples with a rotation higher than $5$ degrees or a translation greater than $1.5$ and $2$ meters for KITTI and DDAD datasets, respectively. We note that this threshold is the only hyperparameter that needs to be tuned in our approach.
IV-C Comparison with monocular depth esimation methods
TABLE I: Quantitative results on the NYUv2 dataset [31] (indoor). (ZS) means Zeros-Shot, (FT) stands for Fine-Tuned on NYUv2.
Methods $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
ZeroDepth [6] (ZS) 0.926 0.081 0.338
Metric3D [29] (ZS) 0.944 0.083 0.310
Metric3D V2 [30] (ZS) 0.975 0.063 0.251
@@ IV-C Comparison with monocular depth esimation methods
Ours w/ LiDAR 1 beam 0.939 0.063 0.652
Ours w/ LiDAR 16 beams 0.976 0.039 0.454
[…]
TABLE II: Quantitative results on the KITTI dataset [35] (outdoor). (ZS) means Zeros-Shot, (FT) stands for Fine-Tuned on KITTI.
Methods $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
ZeroDepth [6] (ZS) 0.910 0.102 4.044
Metric3D [29] (ZS) 0.964 0.058 2.770
Metric3D V2 [30] (ZS) 0.974 0.052 2.511
ZoeDepth [27] (FT) 0.971 0.057 2.281
ZeroDepth [6] (FT) 0.968 0.053 2.087
Metric3D V2 [30] (FT) 0.985 0.044 1.985
ScaleDepth [28] (FT) 0,980 0,048 1,987
@@
Ours w/ LiDAR 1 beam 0.891 0.131 3.096
Ours w/ LiDAR 16 beams 0.967 0.060 2.695
Ours w/ LiDAR 32 beams 0.967 0.060 2.673
Ours w/ SFM (SIFT [37]) 0.893 0.103 3.920
Ours w/ SFM (OmniGlue [38]) 0.925 0.093 3.562
TABLE III: Quantitative results on different zero-shot indoor benchmarks. FT NYUv2 stands for Fine-Tuned on NYUv2 dataset [31].
Methods SUN-RGBD [32] IBIMS-1 [33] DIODE Indoor [34]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
Metric3D [29] 0.144 0.252
@@
TABLE IV: Quantitative results on different zero-shot outdoor benchmarks. FT KITTI stands for Fine-Tuned on KITTI dataset [35].
Method DIODE Outdoor [34] DDAD [36]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
ZoeDepth (FT KITTI) [27] 0.835 0.129 7.108
@@
Depth Anything V1 [2] (FT KITTI) 0.288 0.794 6.641 0.886 0.105 5.931
Our w/ LiDAR 1 beam 0.689 0.880 6.222 0.706 0.326 9.229